OpenStudy (anonymous):

find the 22nd term of the arithmetic sequence 101,105,109,113

OpenStudy (anonymous):

185

OpenStudy (anonymous):

t(n) = 101 + (n-1)4

OpenStudy (anonymous):

all yu freakin do is go up 4 every time..dont be lazy! yu can do that urself!!!1!!

OpenStudy (anonymous):

i did it in 2 seconds

OpenStudy (anonymous):

claqe you are offically banned from answeing any of my questions

OpenStudy (anonymous):

oooo so.! do i care/

OpenStudy (anonymous):

neither do i

OpenStudy (anonymous):

ok then...just stay in school and do nothing! i bet yu are in summer school

OpenStudy (amistre64):

A{0} = ? A{1} = A{0} + 4 = 101 A{2} = A{0} + 4 + 4 = 105 A{3} = A{0} + 4 + 4 + 4 = 109 A{n} = A{0} +4n 101 = A{0} + 4 ; A{0} = 107 A{n} = 107 + 4n .... i did it yay!!!

OpenStudy (amistre64):

......lol almost 101 - 4 = 97 not 107....

OpenStudy (anonymous):

i do do all my work so talk wat you know.. that you don't know DON'T say nun.!

OpenStudy (anonymous):

kudos to you amistre64 for trying :)